# Multifidelity Simulation-based Inference for Computationally Expensive Simulators

### v2

@@ Appendix C Task 2: Multicompartmental single neuron model
We are interested in inferring the densities of two prominent ion channels $\bar{g}_{\mathrm{Na}}$ and $\bar{g}_{\mathrm{K}}$. The low- and high-fidelity models differ in the number of compartments per branch: the low-fidelity model has a single compartment per branch, while the high-fidelity model consists of eight compartments per branch.
All simulations were performed using Jaxley (Deistler et al., 2024) over 120 ms. The injection current is a step current of 0.55mV over 100 ms, with a delay of 10ms. The step size of the simulator is 0.025.
When sampling from the prior distribution over parameters, approximately $0.1\%$ of the respective simulations had clearly unrealistic summary statistics: these simulations were iteratively replaced by random draws from the prior distribution until we collected a desired number of valid simulations.
C.1 Posterior distributions
Posterior distributions over the parameters of the multicompartmental neuron model trained on 10k high-fidelity simulations.
Figure 9: Simulation-based calibration and respective posterior distributions for the multicompartmental neuron model trained on 10k high-fidelity simulations.
C.2 Posterior Predictive Checks
MF-NPE gives similar results as NPE with a training set of high-fidelity simulations up to two orders of magnitude smaller than NPE.
Figure 10: Posterior predictives for the multicompartmental neuron model.
Appendix D Task 3: Spiking network model
D.1 High-fidelity model
@@
with $w_{ij}(t)$ the connection strength between neurons $i$ and $j$ (unitless), $\delta_{k}(t)=\sum\delta(t-t_{k}^{*})$ the spike train of pre-synaptic neuron $k$, where $t_{k}^{*}$ denotes the spike times of neuron $k$, and $\delta$ the Dirac delta. All neurons received input from 5k Poisson neurons, with 5% random connectivity and constant rate $r_{\text{ext}}=10$Hz in each simulation. The recurrent connectivity was instantiated with random sparse connectivity (10%). All recurrent synapses in the network ($E$-to-$E$ and $E$-to-$I$, $I$-to-$E$, $I$-to-$I$) underwent variations of spike-timing dependent plasticity (STDP) (Gerstner & Kistler, 2002; Confavreux et al., 2023). Given the learning rate $\eta$, the weights between the neurons $i$ and $j$ of connection type $X$-to-$Y$ evolved over time as:
@@
$\frac{\mathrm{d}x_{i}}{\mathrm{~{}d}t}=-\frac{x_{i}}{\tau_{\mathrm{XY}}^{\mathrm{pre}}}+\delta_{i}(t)\quad\text{ and }\quad\frac{\mathrm{d}x_{j}}{\mathrm{~{}d}t}=-\frac{x_{j}}{\tau_{\mathrm{XY}}^{\text{post }}}+\delta_{j}(t)$ (11)
@@
D.2 Low-fidelity model
Following previous work (Confavreux et al., 2023; Vogels et al., 2011; Dayan & Abbott, 2001), a (partial) mean-field theory applied to the $E$-to-$E$ and $E$-to-$I$ connections in the model described above gave:
@@ D.2 Low-fidelity model
Appendix E Prior bounds across tasks
@@
Table 2: Prior bounds for i.i.d. Gaussian samples and the Ornstein-Uhlenbeck process.
parameter name lower bound upper bound
$\mu$ 0.1 3.0
$\sigma$ 0.1 0.6
$\gamma$ 0.1 1.0
$\mu_{\mathrm{offset}}$ 0.0 4.0
@@
Table 4: Prior bounds for each synapse type ($E$-to-$E$, $E$-to-$I$, $I$-to-$E$ and $I$-to-$I$) for the spiking neural network and mean-field model.
@@
$\alpha$ $-2$ $2$
$\beta$ $-2$ $2$
$\gamma$ $-2$ $2$
$\kappa$ $-2$ $2$
Appendix F Sequential Algorithms
F.1 MF-TSNPE
@@
  Input: $N$ pairs of $(\boldsymbol{\theta},\boldsymbol{x}_{\text{L}})$; conditional density estimators $q_{\psi}(\boldsymbol{\theta}|\boldsymbol{x}_{\text{L}})$ and $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$ with learnable parameters $\psi$ and $\phi$; early stopping criterion $S$; simulator $p_{\mathrm{H}}(\boldsymbol{x}|\boldsymbol{\theta})$; prior $p(\boldsymbol{\theta})$; number of rounds $R$; $\epsilon$ that defines the highest-probability region ($\mathrm{HPR}_{\epsilon}$); number of high-fidelity simulations per round $M$.
@@
  $\mathcal{L}(\psi)=\frac{1}{N}\sum_{i=1}^{N}-\log q_{\psi}\left(\boldsymbol{\theta}_{i}|\boldsymbol{x}^{\text{L}}_{i}\right)$
@@
        $\mathcal{L}(\phi)=\frac{1}{M}\sum_{i=1}^{M}-\log q_{\phi}\left(\boldsymbol{\boldsymbol{\theta}}^{(r)}_{i}|\boldsymbol{x}^{(r)}_{i}\right)$.
@@
     $\tilde{p}(\boldsymbol{\theta})\propto p(\boldsymbol{\boldsymbol{\theta}})\cdot\mathbbm{1}_{\theta\in\mathrm{HPR}_{\epsilon}}$
@@ F.2 A-MF-TSNPE
  Input: $N$ pairs of $(\boldsymbol{\theta},\boldsymbol{x}_{\text{L}})$; conditional density estimator $q_{\psi}(\boldsymbol{\theta}|\boldsymbol{x}_{\text{L}})$ with learnable parameters $\psi$ and and ensemble of conditional density estimators $\{q^{e}_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})\}^{e}_{E}$, each with independent $\phi$; early stopping criterion $S$; simulator $p_{\mathrm{H}}(\boldsymbol{x}|\boldsymbol{\theta})$; prior $p(\boldsymbol{\theta})$; number of rounds $R$; $\epsilon$ that defines the highest-probability region ($\mathrm{HPR}_{\epsilon}$); number of high-fidelity simulations per round $M$.
@@
  $\mathcal{L}(\psi)=\frac{1}{N}\sum_{i=1}^{N}-\log q_{\psi}\left(\boldsymbol{\theta}_{i}|\boldsymbol{x}^{\text{L}}_{i}\right)$
@@
     $\boldsymbol{\theta}^{(r)}_{\mathrm{prop}}\sim\tilde{p}(\boldsymbol{\theta})$, generate $M-B$ samples from proposal
@@
           $\mathcal{L}(\phi)=\frac{1}{M}\sum_{i=1}^{M}-\log q^{e}_{\phi}\left(\boldsymbol{\boldsymbol{\theta}}^{(r)}_{i}|\boldsymbol{x}^{(r)}_{i}\right)$.
@@
     $\tilde{p}(\boldsymbol{\theta})\propto p(\boldsymbol{\boldsymbol{\theta}})\cdot\mathbbm{1}_{\theta\in\mathrm{HPR}_{\epsilon}}$
@@
All experiments were ran with $R=5$ rounds, $\epsilon=1e^{-6}$, and an ensemble of 5 networks. We set $B=.2M$ to mitigate the concern of biasing the posterior with parameters selected with the acquisition function.

### v3

@@ Appendix G Task 2: Multicompartmental single neuron model
We are interested in inferring the densities of two prominent ion channels $\bar{g}_{\mathrm{Na}}$ and $\bar{g}_{\mathrm{K}}$.
The low- and high-fidelity models differ in the number of compartments per branch: the low-fidelity model has a single compartment per branch, while the high-fidelity model consists of eight compartments per branch.
All simulations were performed using Jaxley (V 0.8.2) (Deistler et al., 2024) over 120 ms. The injection current is a step current of 0.55mV over 100 ms, with a delay of 10ms. The step size of the simulator is 0.025.
When sampling from the prior distribution over parameters, approximately $0.05-0.1\%$ of the respective simulations had clearly unrealistic summary statistics: these simulations were iteratively replaced by random draws from the prior distribution/proposal or active learning list (depending on the algorithm) until we collected a desired number of valid simulations.
G.1 NRMSE evaluation
In addition to the NLTP metric, we demonstrate that the NRMSE metric yields results that support our conclusions.
Figure 13: NRMSE evaluation for the multicompartmental neuron model.
G.2 Simulation-based calibration and Posterior distributions
Figure 14: Simulation-based calibration (left) and respective posterior distributions for NPE and MF-NPE (right) for the multicompartmental neuron model task. MF-NPE is respectively, pretrained on $10^{3},10^{4},10^{5}$ low-fidelity simulations (dubbed as MF-NPE3, MF-NPE4, and MF-NPE5). All models were trained on $10^{3}$ high-fidelity simulations.
G.3 Posterior Predictive Checks
With only 50 high-fidelity simulations, MF-NPE gives similar accuracy to NPE trained on 1000 simulations (Fig. 15), and for a fixed number of 1000 high-fidelity simulations, MF-NPE5 outperforms NPE (Fig. 16).
Figure 15: Posterior predictives for the multicompartmental neuron model with varying number of high-fidelity simulations.
Figure 16: Posterior predictives for the multicompartmental neuron model for a fixed number of high-fidelity simulations.
G.4 Low and high-fidelity traces
We present simulations with the models with 1- and 8-compartments per dendritic branch (low- and high-fidelity models, respectively) to illustrate that the model outputs are different, given the same parameters.
Figure 17: Simulated membrane potential traces of an L5 pyramidal cell (L5PC) model with Jaxley (Deistler et al., 2024). The low- and high-fidelity models are, respectively, a single-compartment model per dendritic branch versus an eight-compartment model per branch.
Appendix H Task 3: Spiking network model
H.1 High-fidelity model
@@
with $w_{ij}(t)$ the connection strength between neurons $i$ and $j$ (unitless), $\delta_{k}(t)=\sum\delta(t-t_{k}^{*})$ the spike train of pre-synaptic neuron $k$, where $t_{k}^{*}$ denotes the spike times of neuron $k$, and $\delta$ the Dirac delta. All neurons received input from 5k Poisson neurons, with 5% random connectivity and constant rate $r_{\text{ext}}=10$Hz in each simulation. The recurrent connectivity was instantiated with random sparse connectivity (10%). All recurrent synapses in the network ($E$-to-$E$ and $E$-to-II, II-to-$E$, II-to-II) underwent variations of spike-timing dependent plasticity (STDP) (Gerstner & Kistler, 2002; Confavreux et al., 2023). Given the learning rate $\eta$, the weights between the neurons $i$ and $j$ of connection type $X$-to-$Y$ evolved over time as:
@@
$\frac{\mathrm{d}x_{i}}{\mathrm{\penalty 10000\ d}t}=-\frac{x_{i}}{\tau_{\mathrm{XY}}^{\mathrm{pre}}}+\delta_{i}(t)\quad\text{ and }\quad\frac{\mathrm{d}x_{j}}{\mathrm{\penalty 10000\ d}t}=-\frac{x_{j}}{\tau_{\mathrm{XY}}^{\text{post }}}+\delta_{j}(t)$ (14)
@@
H.2 Low-fidelity model
Following previous work (Confavreux et al., 2023; Vogels et al., 2011; Dayan & Abbott, 2001), a (partial) mean-field theory applied to the $E$-to-$E$ and $E$-to-II connections in the model described above gave:
@@ H.2 Low-fidelity model
H.3 Discussion on alternative solutions
We consider the following strategies:
  • pretraining on solely low-fidelity simulations,
  • pretraining on the joint of low- and high-fidelity simulations.
H.3.1 Pretraining on low fidelity samples
This approach follows the main discussion in Sec. 3.1.1, and has also been the main method employed in the paper. We purposefully do not freeze the weights after transfer, allowing the network to retain the flexibility to adapt to high-fidelity simulations.
H.3.2 Pretraining on the joint of LF and HF samples
We examined whether pretraining on the joint distribution of low- and high-fidelity simulations could provide a better initialization for subsequent fine tuning. As shown in Fig. 18, this strategy yields no significant improvement on the first two benchmarking tasks compared to standard MF-NPE. However, we encourage further work to investigate additional variations on this approach to improve the domain adaptation (e.g., domain adaptation through MMD Elsemüller et al. (2025), importance weighting for extremely unbalanced datasets, adversarial discriminative domain adaptation, training a single multifidelity inference network).
Figure 18: MF-(TS)NPE (joint) has been pretrained on both low- and high-fidelity samples.
Appendix I Prior bounds across neuroscience tasks
@@
Table 2: Prior bounds for each synapse type ($E$-to-$E$, $E$-to-II, II-to-$E$ and II-to-II) for the spiking neural network and mean-field model.
@@
$\alpha$ 2-2 22
$\beta$ 2-2 22
$\gamma$ 2-2 22
$\kappa$ 2-2 22
Appendix J Distance between the LF and HF posterior
Both the low and high-fidelity posterior distributions have been trained on $10^{5}$ simulations and evaluated over 10 true observations. In the table below, we focus on cases with two fidelities and measure the distance between the low and high-fidelity models with the MMD and C2ST metrics. We observe that the distance between the posterior distributions is not a direct measure of success in transfer learning. For instance, the posterior distributions of the low-and high-fidelity models of the L5PC neuron are significantly different. However, the network still manages to leverage information between the two simulators (Figure 3), supporting the theoretical results of Tahir et al. (2024).
Transfer learning seems to work less well on the OU process task when the dimensionality of the parameters differs between the low- and high-fidelity models (see Sec. 8). This is observed despite the fact that the distance between the low and high-fidelity posteriors is lower for the OU4 case than for the OU2 case, as the low-fidelity OU2 posterior is highly biased (Fig. 19).
Table 3: Distance between low- and high-fidelity posterior (mean $\pm$ std) for different tasks.
Task MMD C2ST
SLCP $0.13\pm 0.05$ $0.91\pm 0.03$
SIR $0.04\pm 0.03$ $0.57\pm 0.03$
OU2 $1.00\pm 0.11$ $0.98\pm 0.02$
OU3 $0.69\pm 0.087$ $0.98\pm 0.01$
OU4 $0.24\pm 0.05$ $0.90\pm 0.04$
L5PC $0.76\pm 0.23$ $0.99\pm 0.00$
SynapticPlasticity $0.01\pm 0.00$ $0.70\pm 0.02$
J.1 Pairplots
[…]
Appendix K Simulation versus training cost
We tracked the wall-clock run-time for training and simulation stages of the neural density estimator. Computations were performed on nodes each equipped with 4× Intel Xeon Gold 6448H CPUs (32 cores per socket, 128 physical cores, 256 logical CPUs) and approximately 2TB RAM, running Linux 5.14.0. We compare the costs in regimes where the performance of NPE is similar to MF-NPE and A-MF-TSNPE (Fig. 3). Details about the network architecture and hyperparameters are described in Appendix D.1. In cases where many samples had to be generated for active non-amortized schemes (e.g., $10^{5}$ HF samples for the L5PC task; Figure 3), we used multiprocessing over CPUs. The simulations for the third task were parallelized over 913 CPUs.
Table 4: Comparison of methods for the real-world tasks in terms of the number of simulations and computational cost. Total training cost is reported as mean ± standard deviation over 5 network runs.
method # simulations CPU (seconds)
LF HF tot. cost (sim.) tot. cost (train) total cost
L5PC NPE NA $10^{4}$ 4940 70.39 ± 18.32 5010.39 ± 18.32
MF-NPE $10^{4}$ $10^{3}$ 1032 96.94 ± 15.19 1128.94 ± 15.19
A-MF-TSNPE $10^{4}$ 50 607 557.44 ± 52.5 1164.44 ± 52.5
Network NPE NA $10^{4}$ $3\times 10^{6}$ 120.43 3,000,120
MF-NPE $10^{5}$ $10^{3}$ $3\times 10^{5}$ 94.54 300,094
Table 5: Comparison of methods across models in terms of the number of simulations and accuracy. Evaluated using the NLTP metric.
Method # Simulations
Accuracy
(C2ST/NLTP)
LF HF
L5PC NPE NA $10^{4}$ -5.87 ± 0.04
MF-NPE $10^{4}$ $10^{3}$ -5.73 ± 0.05
A-MF-TSNPE $10^{4}$ 50 -5.08 ± 0.27
Network NPE NA $10^{4}$ -4.72 ± 0.01
MF-NPE $10^{5}$ $10^{3}$ -4.08 ± 0.01
Table LABEL:table:train-sim-time shows that the multifidelity approaches make sense when the training cost is significantly lower than the simulation cost, such as in the L5PC and the spiking network model. For instance, in the spiking network task, a single high-fidelity simulation requires approximately 5 CPU minutes, whereas a low-fidelity simulation takes only 0.0008 seconds.
Appendix L MF-NPE for multiple lower-fidelity simulators
Algorithm 2 MF-NPE with multiple fidelities
Input: Simulations $\{(\boldsymbol{\theta},\boldsymbol{x}^{(f)})\}^{F}_{f=1}$ over $F$ fidelities; Early stopping criterion $S$; conditional density estimators $\{q_{\psi}^{(f)}(\boldsymbol{\theta}|\boldsymbol{x}^{(f)})\}^{F}_{f=1}$ with features $\psi$.
for $f=1$ to $F$ do
  $\mathcal{L}(\psi^{(f)})=\frac{1}{N^{(f)}}\sum_{i=1}^{N^{(f)}}-\log q^{(f)}_{\psi}\left(\boldsymbol{\theta}_{i}|\boldsymbol{x}^{(f)}_{i}\right)$ .
  $\mathrm{opt}^{(f)}\leftarrow\mathrm{Adam}(\cdot)$
  if $f>1$ then
   Initialize $q_{\psi}^{(f)}$ with features of trained $q_{\psi}^{(f-1)}$.
  end if
  for epoch in epochs do
   train $q_{\psi}^{(f)}$ to minimize $\mathcal{L}(\psi^{(f)})$ until $S$ is reached.
  end for
end for
Appendix M Sequential Algorithms
M.1 MF-TSNPE
@@
Input: NN pairs of $(\boldsymbol{\theta},\boldsymbol{x}_{\text{L}})$; conditional density estimators $q_{\psi}(\boldsymbol{\theta}|\boldsymbol{x}_{\text{L}})$ and $q_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})$ with learnable parameters $\psi$ and $\phi$; early stopping criterion $S$; simulator $p_{\mathrm{H}}(\boldsymbol{x}|\boldsymbol{\theta})$; prior $p(\boldsymbol{\theta})$; number of rounds $R$; $\epsilon$ that defines the highest-probability region ($\mathrm{HPR}_{\epsilon}$); number of high-fidelity simulations per round $M$.
@@
$\mathcal{L}(\psi)=\frac{1}{N}\sum_{i=1}^{N}-\log q_{\psi}\left(\boldsymbol{\theta}_{i}|\boldsymbol{x}^{\text{L}}_{i}\right)$ .
@@
   $\mathcal{L}(\phi)=\frac{1}{M}\sum_{i=1}^{M}-\log q_{\phi}\left(\boldsymbol{\theta}^{(r)}_{i}|\boldsymbol{x}^{(r)}_{i}\right)$.
@@
  $\tilde{p}(\boldsymbol{\theta})\propto p(\boldsymbol{\theta})\cdot\mathbbm{1}_{\theta\in\mathrm{HPR}_{\epsilon}}$
@@
Input: NN pairs of $(\boldsymbol{\theta},\boldsymbol{x}_{\text{L}})$; conditional density estimator $q_{\psi}(\boldsymbol{\theta}|\boldsymbol{x}_{\text{L}})$ with learnable parameters $\psi$ and and ensemble of conditional density estimators $\{q^{e}_{\phi}(\boldsymbol{\theta}|\boldsymbol{x})\}^{e}_{E}$, each with independent $\phi$; early stopping criterion $S$; simulator $p_{\mathrm{H}}(\boldsymbol{x}|\boldsymbol{\theta})$; prior $p(\boldsymbol{\theta})$; number of rounds $R$; $\epsilon$ that defines the highest-probability region ($\mathrm{HPR}_{\epsilon}$); number of high-fidelity simulations per round $M$.
@@
$\mathcal{L}(\psi)=\frac{1}{N}\sum_{i=1}^{N}-\log q_{\psi}\left(\boldsymbol{\theta}_{i}|\boldsymbol{x}^{\text{L}}_{i}\right)$ .
@@
  $\boldsymbol{\theta}^{(r)}_{\mathrm{prop}}\sim\tilde{p}(\boldsymbol{\theta})$, generate MBM-B samples from proposal
@@
    $\mathcal{L}(\phi)=\frac{1}{M}\sum_{i=1}^{M}-\log q^{e}_{\phi}\left(\boldsymbol{\theta}^{(r)}_{i}|\boldsymbol{x}^{(r)}_{i}\right)$.
@@
  $\tilde{p}(\boldsymbol{\theta})\propto p(\boldsymbol{\theta})\cdot\mathbbm{1}_{\theta\in\mathrm{HPR}_{\epsilon}}$
@@
All experiments were ran with $R=5$ rounds, $\epsilon=1e^{-6}$, and an ensemble of 5 networks. The addition of an acquisition function biases the proposal distribution, causing the density estimate to diverge from the true posterior. In principle, this could be addressed by using atomic proposals (Greenberg et al., 2019), but given that such an approach suffers from posterior leakage, we do not introduce a proposal correction in order to retain the well-behaved loss function in TNSPE. We argue that the benefit of informative samples would outweigh the potential bias, as long as the percentage of samples selected from the acquisition function would be small compared to the proposal samples. Therefore, we set $B=.2M$ to mitigate the concern of biasing the posterior with parameters selected with the acquisition function.